OpenStudy (anonymous):

a merry-go-round ride has outer horse figures that go up and downw exactly three times during one rotation of the carousel. One of the high-points for each horse occurs where a rider is just close enough to reach out and try to grab a metal ring from a mechanical dispenser. If the rider succeeds in grabbing a brass ring instead of an iron ring, the rider has won a free ride on the merry-go-round. Hence the expression, "reaching for the brass ring." Write a sinusoidal function that models the height of one of the horse figures as a function of the rotation of the main carousel with beta=0 at th

OpenStudy (anonymous):

ring dispenser. The amplitude of the horse's vertical motion is 1.1 meters aroudn the average height of 1.6. answer choices are: y= 1.1sin(3beta+pi/2)+1.6 y=.505sin(3/2beta+pi/2)+1.6 y=.505sin(2/3beta-pi/2)+1.6 y=1.1sin(1/3beta-pi/2)+1.6

OpenStudy (anonymous):

whichever has k = 2pi/3

OpenStudy (anonymous):

since period = 2pi/k

OpenStudy (anonymous):

which one would tht be?